OpenStudy (anonymous):

Simplify 10a2b(a2 - 3a - 4)

OpenStudy (waleed_imtiaz):

=10a^2(b)(a^2-3a-4) =10a^2(b)[a^2-4a+a-4] =10a^2(b)[a(a-4)+1(a-4)] =10a^2(b)(a+1)(a-4)

OpenStudy (anonymous):

THANKS SOO MUCH!!

OpenStudy (anonymous):

CAN U answer 1 more for me??

OpenStudy (waleed_imtiaz):

Mm... okay. Let me see.... WIll try......

OpenStudy (anonymous):

Multiply and simplify: (3p - 8)(4p - 2) thanswer choices are: 10a4b - 30a2b - 4 10a4b - 30a3b - 40a2b 10a4b - 70a2b 10a4b - 30a3b - 4

OpenStudy (waleed_imtiaz):

what are U saying man.... See ur choices and the question difference.... U are giving p in question and in choices, no p ?why ?

OpenStudy (anonymous):

sorry meant: 12p2 - 26p - 16 12p2 - 38p + 16 12p2 - 16 12p2 + 16d

OpenStudy (waleed_imtiaz):

It should be 12p2-38p+16.... simple multiplication

OpenStudy (waleed_imtiaz):

U got it ?